# CMAR-Net: Accurate Cross-Modal 3D SAR Reconstruction of Vehicle Targets with Sparse Multi-Baseline Data

### v3

@@ Abstract
Multi-baseline Synthetic Aperture Radar (SAR) three-dimensional (3D) tomography is a crucial remote sensing technique that provides 3D resolution unavailable in conventional SAR imaging. However, achieving high-quality imaging typically requires multi-angle or full-aperture data, resulting in significant imaging costs. Recent advancements in sparse 3D SAR, which rely on data from limited apertures, have gained attention as a cost-effective alternative. Notably, deep learning techniques have markedly enhanced the imaging quality of sparse 3D SAR. Despite these advancements, existing methods primarily depend on high-resolution radar images for supervising the training of deep neural networks (DNNs). This exclusive dependence on single-modal data prevents the introduction of complementary information from other data sources, limiting further improvements in imaging performance. In this paper, we introduce a Cross-Modal 3D-SAR Reconstruction Network (CMAR-Net) to enhance 3D SAR imaging by integrating heterogeneous information. Leveraging cross-modal supervision from 2D optical images and error transfer guaranteed by differentiable rendering, CMAR-Net achieves efficient training and reconstructs highly sparse multi-baseline SAR data into visually structured and accurate 3D images, particularly for vehicle targets. Extensive experiments on simulated and real-world datasets demonstrate that CMAR-Net significantly outperforms state-of-the-art sparse reconstruction algorithms based on compressed sensing (CS) and deep learning (DL), with average improvements of 75.83% in PSNR and 47.85% in SSIM. Furthermore, our method eliminates the need for time-consuming full-aperture data preprocessing and relies solely on computer-rendered optical images, significantly reducing dataset construction costs. This work highlights the potential of deep learning for multi-baseline SAR 3D imaging and introduces a novel framework for radar imaging research through cross-modal learning.
keywords:
@@
Multi-baseline SAR , Sparse imaging , 3D reconstruction , Cross-modal learning , Information fusion
[…]
[bit-phy2]organization=Beijing Key Laboratory of Nanophotonics and Ultrafine Optoelectronic Systems, School of Physics, Beijing Institute of Technology,addressline=No.5 Zhongguancun South Street, city=Haidian District, postcode=100081, state=Beijing, country=China
1 Introduction
Multi-baseline SAR 3D tomography is a highly promising technique for SAR-based 3D reconstruction, due to its ability to provide complete threedimensional spatial resolution[1]. Multi-baseline SAR forms a synthetic aperture in the elevation dimension by performing multiple flights at varying altitudes, providing height resolution, enabling 3D imaging. This technology has significant potential in applications such as high-precision geographic remote sensing, urban 3D mapping, and detailed target interpretation [2]. In contrast, traditional 2D SAR generates a two-dimensional projection of a 3D target in the slant-range azimuth plane, leading to distortions such as overlap and perspective shortening, which result in poor image clarity and limited interpretability [3]. SAR 3D imaging, on the other hand, captures the spatial distribution of scattering centers in three dimensions, providing an accurate and intuitive representation of the target’s structure. As a result, it has gradually become a key focus of SAR technology research.
However, applying multi-baseline SAR for 3D observation presents several challenges in practical scenarios. To obtain a 3D image, multi-baseline SAR requires observations from multiple long baselines or even circular trajectories around the target area to achieve high angular resolution[4, 5, 6]. Due to constraints such as terrain, weather, and flight conditions, observations are often limited to a few angles or apertures along restricted flight paths, making it difficult to meet the data requirements. Moreover, the large volume of data generated by multi-baseline observations places significant demands on data computation and storage [7]. To address these issues, researchers have focused on developing sparse imaging algorithms that can reconstruct detailed 3D target images from limited and incomplete data. Existing methods can generally be classified into two main categories: those based on compressed sensing and those based on deep learning.
Compressed sensing (CS) has long been the mainstream method for reconstructing signals from incomplete and sparse measurements, and it was applied early in multi-baseline SAR 3D imaging [8, 9]. These methods treat the sparse multi-baseline SAR 3D imaging problem as a signal recovery model. By introducing various penalty functions and optimization techniques, they are able to reconstruct high-resolution target images from incomplete measurements. Typical works, such as [9, 10], have reduced the reliance of multi-baseline SAR 3D imaging on complete measurements. However, CS-based methods are computationally intensive, especially when dealing with 3D data. Additionally, CS solvers are highly sensitive to parameter settings, which limits their generalization and hampers their further development and application[11, 12].
Deep learning (DL) techniques have been increasingly applied in the field of sparse multi-baseline SAR 3D imaging in recent years. Researchers have explored deep neural networks (DNNs) to learn image priors from training data and apply them for sparse imaging[13, 14, 15, 16, 17]. Thanks to their parallel structure and inference-based (non-iterative) mode, DL methods offer more efficient and stable imaging results compared to CS-based approaches[18]. Based on their implementation, DL-based methods can be divided into two types.
@@ 1 Introduction
The first type of method embeds NN models into the CS framework, replacing the traditional iterative process with cascaded DNNs that model the nonlinear optimization components. Notable works include [13, 19, 20, 21]. After extensive training, the network learns data priors and optimization parameters, eliminating the need for manual configuration. Unlike CS-based methods, this approach requires only a single inference to achieve reconstruction accuracy comparable to CS algorithms, significantly reducing the computational complexity.
@@
The second type of method combines traditional imaging algorithms with DNNs. It first applies traditional algorithms to pre-imaging sparse data, and then uses DNNs to refine the results. For example, in [22], a 3D U-Net was used to improve back-projection (BP) pre-imaging, enabling high-resolution reconstruction of distorted BP images. Similarly, in [23], a GAN-based sparse-aspects-completion network (SACNet) enhanced the results of CS pre-imaging, demonstrating that a network trained on simulated data can still achieve good target reconstruction on real-world data. These methods frame the improvement of pre-imaging results as an image enhancement or denoising task. Thanks to the strong data representation ability of DNNs and extensive training, they can quickly and stably reconstruct high-resolution 3D target images with a single network inference, even under conditions of highly sparse observation data. As a result, these methods have become the state-of-the-art (SOTA) for sparse SAR 3D reconstruction.
@@
In summary, DL-based sparse multi-baseline SAR 3D reconstruction methods demonstrate significant potential in enhancing both imaging efficiency and quality[18]. However, current learning-based approaches still face several significant practical challenges.
@@
    DL-based methods learn the mapping between training data pairs, which are derived from a single modality (electromagnetic information). This constraint on resolution, imposed by the electromagnetic nature, limits further improvements in image quality.
@@
    As a data-driven technology, DL-based methods rely heavily on the quality of the training data. Since high-resolution 3D SAR images are required for supervision, existing methods often necessitate time-consuming preprocessing of full-aperture data. This process consumes significant storage space and demands high-performance real-time processing platforms, making it costly and limiting the practical deployment of such algorithms.
@@
    The pre-imaging results fed into the network for high-resolution reconstruction are sensitive to the observation geometry of multi-baseline SAR, including the varying scattering characteristics of anisotropic targets at different angles, and are highly susceptible to noise interference. As a result, the reconstruction network must have strong generalization capabilities.
To overcome the aforementioned limitations, particularly the resolution constraints caused by single-modality information, in this paper, we introduce cross-modal learning. The cross-modal learning explores the intrinsic relationships between different data modalities, seeks shared representations, and integrates heterogeneous information. This concept has recently been successfully applied to problems such as SAR image translation[24, 25, 26].
Here, we propose a sparse multi-baseline SAR 3D reconstruction method based on cross-modal supervision and design a CMAR-Net to significantly enhance the quality of 3D SAR images. The supervisory information for training CMAR-Net is directly derived from aligned optical 2D images. The primary challenge of using 2D images to supervise a 3D imaging network lies in the difference in data formats. To address this, we employ differentiable rendering techniques to compute the consistency error between 2D pixels and 3D voxels, ensuring that this error can be back-propagated through the network. The main contributions of this paper are as follows:
  1. 1.
    To the best of our knowledge, this work is the first to introduce the concept of cross-modal learning into SAR 3D reconstruction. This approach enables high-resolution reconstructions with complete visual structure and distinct features, overcoming the resolution constraints inherent to electromagnetic images.
  2. 2.
    We propose a novel CMAR-Net that integrates a unique data augmentation strategy and a projection-reprojection module, enhancing both robustness and generalization. The network is trained on a limited set of simulated data and successfully validated on real-world data without any fine-tuning.
  3. 3.
    The proposed method only requires 2D optical images as supervisory data, eliminating the need for high-resolution images derived from full-aperture data. This simplifies the construction of high-quality imaging datasets, facilitating the broader application of 3D SAR reconstruction technology.
@@
    Extensive experiments demonstrate that our method achieves outstanding 3D target reconstruction results even under low signal-to-noise ratio (SNR) conditions and with highly sparse angular measurements. Image quality improves 75.83% in PSNR and 47.85% in SSIM compare to SOTA methods.
2 Methodology
2.1 Method framework
The proposed multi-baseline SAR 3D reconstruction method using sparse data is shown in Figure 1. Our method includes two modules: the pre-imaging module and the cross-modal reconstruction module. The pre-imaging module processes sub-aperture data and combines them noncoherently to produce a coarse, incomplete 3D image. The cross-modal reconstruction module, which includes a DNN trained with optical image supervision, processes the pre-imaging result to generate realistic 3D image with clear contours and complete structures. The details of each module are explained below.
@@ 2.1 Method framework
Figure 1: The framework of the proposed sparse multi-baseline SAR cross-modal reconstruction method.
2.2 Pre-imaging module
Figure 1(left) illustrates the multi-baseline SAR multi-angle observation geometry. The radar platform performs multi-angle measurements around the target, shown as colored line segments. At each sub-angle, multiple straight-line flights at different altitudes generate synthetic apertures in both the azimuth and elevation dimensions, represented by dot clusters. Under far-field conditions, frequency-domain echo data are collected at different azimuth angles and heights for each sub-aperture, forming a 3D circular k-space observation of the target. After interpolation and coordinate transformation, the spatial positions of the target’s scattering centers can be obtained through a 3D inverse Fourier transform[8].
In the pre-imaging stage, phase errors from trajectory inaccuracies are corrected first, and the frequency-domain data in polar coordinates are then interpolated into Euler coordinates. A 3D inverse Fourier transform then reconstructs the target’s 3D image for each sub-aperture. Finally, the images from all sub-apertures are non-coherently summed to generate the pre-imaging result, which serves as input to the cross-modal reconstruction module.
2.3 Cross-modal reconstruction module
The cross-modal reconstruction module takes pre-imaging result as input and uses CMAR-Net to generate high-resolution 3D images that accurately represent the target’s physical shape. As shown in Figure 2, the network follows an encoder-decoder architecture with skip connections. The encoder extracts structural features from the incomplete pre-imaging result, while the decoder reconstructs the vehicle’s 3D image from the low-dimensional latent representation. To further improve feature representation, a projection-reprojection (PRP) unit is added near the bottleneck layer. At the output, a differentiable rendering module is introduced to enable cross-modal supervision. The following sections describe each module in detail.
2.3.1 Network architecture
@@
Figure 2 illustrates the detailed structure and training process of the CMAR-Net. The network consists of a contraction path (on the left), an expansion path (on the right), and a PRP unit. The contraction path includes four downsampling layers, each consisting of a 3D convolutional layer followed by a max-pooling layer. The convolutional layers apply LeakyReLU activation and progressively increase the number of feature channels, while the pooling layers reduce the data dimensionality. The contraction path ends with a connection to the PRP unit.
@@
The expansion path mirrors the contraction path, consisting of four upsampling layers and a single convolutional layer. Each upsampling layer uses a 3D transposed convolution with ReLU activation. Feature maps from the expansion path are combined with those from the contraction path at the same resolution through skip connections before being passed through the upsampling layers. These layers reduce the number of feature channels while increasing spatial resolution. The final upsampling layer produces a 64-channel output, then refined by a output convolution layer to smooth and integrate information across channels, ensuring a more refined reconstruction.
@@
2.3.2 Projection-Reprojection unit
We design a PRP unit between the encoder and decoder to improve representation and generalization.
The structure of the PRP unit is detailed in Figure 3. The projection part consists of two fully connected layers. The first fully connected layer, activated by LeakyReLU, compresses the input feature map into a low-dimensional vector. The second fully connected layer eliminates the activation and further compresses the feature’s dimensionality to produce the latent representation $z$. The reprojection part reverses the projection process by symmetrically expanding the feature vector and reshaping it back to the original feature map size. The only difference is that the final fully connected layer uses ReLU activation to ensure consistency in the data dynamics within the network.
@@ 2.3.2 Projection-Reprojection unit
2.4 Differentiable volume rendering
Optical images contain rich, intuitive structural information, and, due to their passive imaging mechanism, high-resolution optical images are often easier to obtain in practical applications compared to radar electromagnetic images. However, integrating the advantages of 2D optical images into 3D SAR imaging requires bridging the dimensional gap between the two data types.
Differentiable volume rendering offers a promising solution. This technique converts 3D volumetric data into 2D images, and its differentiable nature allows gradient-based optimization of 3D structures using a 2D loss function, facilitating effective supervision of 3D reconstruction through 2D images. In recent years, it has been widely used in computational graphics[27], novel view synthesis[28] and point cloud completion[29, 30]. In our network, we incorporate a differentiable volume rendering module at the end of the CMAR-Net to generate 2D images from multiple views. By supervising the network with optical images, we leverage cross-modal benefits to improve the quality of the 3D SAR reconstruction.

### v4

# CMAR-Net: Accurate Cross-Modal 3D SAR Reconstruction of Vehicle Targets with Sparse-Aspect Multi-Baseline Data

@@ Abstract
Sparse-aspect multi-baseline Synthetic Aperture Radar (SAR) three-dimensional (3D) tomography is a crucial remote sensing technique. Compared to full-aspect observation, it needs only a few observation aspects to achieve a sufficiently clear 3D scene reconstruction, providing a cost-effective alternative. In the past, compressive sensing (CS) was the mainstream approach for sparse 3D SAR imaging. Recently, deep learning (DL) revolutionizes this field through its powerful data-driven representation capabilities and efficient inference characteristics. However, existing DL methods primarily depend on high-resolution radar images for supervising the training of deep neural networks (DNNs). This unimodal approach precludes the incorporation of complementary information from other data sources, thereby limiting potential improvements in imaging performance. In this paper, we propose a Cross-Modal 3D-SAR Reconstruction Network (CMAR-Net) that enhances 3D SAR imaging by fusing heterogeneous information. Leveraging cross-modal supervision from 2D optical images and error transfer guaranteed by differentiable rendering, CMAR-Net achieves efficient training and reconstructs highly sparse-aspect multi-baseline SAR image into visually structured and accurate 3D images, particularly for vehicle targets. Extensive experiments on simulated and real-world datasets demonstrate that CMAR-Net significantly outperforms state-of-the-art sparse reconstruction algorithms based on CS and DL, with average improvements of 75.83% in PSNR and 47.85% in SSIM. Furthermore, our method eliminates the need for time-consuming full-aperture data preprocessing and relies solely on computer-rendered optical images, significantly reducing dataset construction costs. This work highlights the potential of cross-modal learning for multi-baseline SAR 3D imaging and introduces a novel framework for radar imaging research.
Index Terms:
@@
Multi-baseline SAR, Sparse imaging, 3D reconstruction, Cross-modal learning, Information fusion.
I Introduction
Sparse-aspect multi-baseline SAR is a highly promising SAR-based 3D reconstruction technique. On one hand, it retains the essential capabilities of conventional 3D SAR imaging, such as capturing the spatial distribution of scatterers and providing an accurate and intuitive representation of target structures[1, 2]. On the other hand, sparse-aspect multi-baseline SAR achieves relatively high-quality volumetric reconstructions using only a limited number of observation aspects, thereby relaxing the stringent sampling requirements often imposed in practical applications and significantly reducing the overall imaging cost[3]. Owing to these advantages, sparse-aspect multi-baseline SAR holds significant potential for a wide range of applications, including high-precision geospatial remote sensing[4], urban 3D mapping[5], and detailed target interpretation[6].
In typical sparse multi-baseline SAR application scenarios, scatterers are often sparsely distributed along the elevation dimension, with only a few dominant scatterers located within each range-azimuth resolution cell[7]. As a result, CS-based methods have long been the dominant approach for addressing the 3D SAR inversion problem[8, 9, 10]. These methods formulate the sparse imaging problem as a signal recovery model. By incorporating various penalty functions and optimization techniques, they can reconstruct high-resolution target images from incomplete measurements. However, CS-based approaches also suffer from several limitations. The iterative optimization process is computationally intensive[11, 12], and in 3D scenarios, the computational complexity grows exponentially with data dimensionality, rendering large-scale processing increasingly intractable. Moreover, CS solvers are inherently sensitive to noise, require careful hyperparameter tuning, and often lack robustness and generalization capability across diverse imaging conditions[13].
With the rapid advancement of DL technologies and their widespread adoption across various domains, DL-based sparse reconstruction methods have emerged as a promising solution to address the aforementioned limitations. DL is a data-driven approach that learns the underlying features of a scene directly from training data. By employing hierarchical model architectures, it captures complex nonlinear mappings between inputs and outputs. In recent years, numerous studies have explored the use of deep neural networks (DNNs) for tackling sparse imaging tasks[14, 15, 16, 17, 18]. These methods can generally be categorized into two main types.
@@ I Introduction
The first type embeds DNNs into the CS framework by replacing the traditional iterative optimization process with cascaded DNN modules that approximate its nonlinear components. These approaches are commonly referred to as unfolding methods[19, 20, 21, 22, 14]. Through extensive training, these networks learn data priors and optimization parameters, eliminating the need for manual configuration. Owing to their parallel architecture and inference-based (i.e., non-iterative) processing paradigm, these approaches require only a single forward pass to achieve reconstruction accuracy comparable to that of traditional CS algorithms, thereby significantly reducing computational complexity.
@@
The second type approach employs a sequential architecture, cascading DNNs with traditional imaging algorithms in a ”two-step” process. First, a conventional algorithm produces an initial 3D SAR image. This image is then refined by a DNN to enhance the imaging quality. For example, in [23], a 3D U-Net was utilized to improve back-projection (BP) results, enabling high-resolution reconstruction from distorted BP images. Similarly, in [3], a GAN-based sparse-aspects-completion network (SACNet) enhanced the results of CS results, demonstrating that a network trained on simulated data can still achieve effective target reconstruction on real-world data. These methods treat the improvement of 3D SAR imaging results as an image enhancement or denoising task. Leveraging the powerful data representation capabilities of DNNs and extensive training, they can rapidly and reliably reconstruct high-resolution 3D target images, even under conditions of highly sparse observation data.
@@
In summary, DL-based sparse-aspect multi-baseline SAR 3D reconstruction methods demonstrate significant potential in enhancing both imaging efficiency and quality[13]. However, current learning-based approaches still face significant practical problems.
@@
    DL-based methods learn the mapping relationships between training data pairs, which are derived from the single modality (electromagnetic information). This reliance on electromagnetic constraints imposes inherent limits on the achievable resolution and restricts further improvements in image quality.
@@
    As a data-driven technology, the performance of DL-based methods heavily depends on the quality of training data. Most existing frameworks utilize high-resolution 3D SAR images as supervision, which necessitates imaging preprocessing of full-aperture data prior to training. This process demands substantial computational time and storage space, or high-performance real-time processing platforms.
To address the aforementioned limitations—particularly the resolution constraints imposed by single-modality data—we introduce cross-modal learning in this work. The cross-modal learning explores the intrinsic relationships between different data modalities, seeks shared representations, and integrates heterogeneous information. This concept has recently been successfully applied to problems such as cross-modal retrieval[24], image captioning[25], and particularly SAR-to-optical image translation in remote sensing[26, 27, 28].
Here, we propose a sparse-aspect multi-baseline SAR 3D reconstruction framework driven by cross-modal learning, implemented via a novel CMAR-Net, which significantly improves 3D SAR imaging quality. Unlike conventional methods requiring 3D supervision, our network is trained using aligned 2D optical images as supervisory signals. The key challenge lies in bridging the dimensional disparity between 2D pixels and 3D voxels. To resolve this, we adopt differentiable rendering to compute consistency errors that are back-propagatable through the network.
The main contributions of this work are summarized as follows:
  1. 1.
    First cross-modal learning framework for 3D SAR reconstruction: To our knowledge, this is the first application of cross-modal learning to 3D SAR reconstruction. Our method achieves high-resolution reconstructions with complete visual structures and sharp features, overcoming the inherent resolution limitations of electromagnetic imaging.
  2. 2.
    Robust CMAR-Net architecture: We design a novel network incorporating a data augmentation strategy and a projection-reprojection module to enhance robustness and generalization. Remarkably, CMAR-Net trained on simulated data achieves strong performance on real-world datasets without any fine-tuning.
@@
    Reduced supervision requirements: The proposed method only necessitates 2D optical images as supervision, eliminating dependency on high-resolution full-aperture data. This significantly simplifies the creation of high-quality imaging datasets, broadening the practicality of 3D SAR reconstruction.
@@
    State-of-the-art performance: Extensive experiments show our method excels under low SNR and highly sparse aspect, outperforming existing approaches by 75.83% in PSNR and 47.85% in SSIM.
This paper is organized as follows: Section II introduces the imaging geometry, overall framework of the proposed method, and module details. Section III presents the experimental setup, results, and analysis. Section IV concludes the work.
II Methodology
Figure 1: Multi-aspect, multi-baseline SAR observation geometry.
@@ II Methodology
Figure 2: The framework of the proposed sparse-aspect multi-baseline SAR cross-modal reconstruction method.
II-A Multi-aspect, Multi-baseline SAR Observation Geometry
The geometry of multi-aspect, multi-baseline SAR imaging is illustrated in Figure 1. In each observation aspect, the radar platform follows a straight flight path at different altitudes, forming observation apertures in both the azimuth and height dimensions, represented by evenly spaced points in the figure. To capture anisotropic scattering information and achieve precise 3D reconstruction of the target’s structure, the radar platform must conduct multi-aspect observations around the target, measuring the target’s scattering characteristics from various directions, with different colors used in the figure to differentiate the observation aspects. However, in practical applications, the available observation aspects are often sparse due to limitations in flight conditions, acquisition costs, or platform configurations.
II-B Method framework
The framework of the proposed sparse-aspect multi-baseline SAR 3D reconstruction method is illustrated in Figure 2. It consists of two main processing modules: the multi-aspect 3D SAR imaging module and the cross-modal reconstruction module. First, classical 3D SAR imaging process is performed on the sparse-aspect, multi-baseline SAR data. This step includes sub-aspect imaging and non-coherent addition of multi-aspect images. Subsequently, the proposed cross-modal reconstruction network, CMAR-Net, further enhances the 3D SAR image. As shown in the figure, using a vehicle target as an example, CMAR-Net directly takes the 3D SAR image as input and produces a clearer and more structurally complete 3D reconstruction. During training, optical images of vehicles—spatially aligned with the SAR images—are introduced as supervision signals. This allows for quantitative evaluation of the 3D reconstruction accuracy and establishes a cross-modal supervision mechanism, thereby enabling modality transfer from SAR to optical. The details of each module are explained below.
II-C Multi-aspect 3D SAR imaging module
The detailed process of multi-aspect 3D SAR imaging is illustrated in the upper part of Figure 2. The target is observed from several sub-aspect angles, and we begin by independently imaging each sub-aspect. For each sub-aspect, after motion compensation and phase calibration, the polar-format frequency domain data are interpolated onto a Cartesian grid, and a three-dimensional inverse Fourier transform is then applied to reconstruct the target’s spatial-domain image[8]. Finally, all sub-aspect images are non-coherent added to integrate multi-aspect scattering information and generate the final 3D SAR image of the target[29].
II-D Cross-modal reconstruction module
The cross-modal reconstruction module consists of a cross-modal reconstruction network, CMAR-Net, as shown in the lower part of Figure 2. It takes 3D SAR image data as input and utilizes spatially aligned multi-view 2D optical images as supervision signals. During training, the network minimizes the differences between the projected views of the 3D reconstruction result and the corresponding reference optical images from multiple viewpoints, encouraging the reconstructed 3D geometry to align accurately with the structures observed in the optical domain. The detailed architecture and training process of CMAR-Net are depicted in Figure 3. The following sections describe each component in detail.
II-D1 Network architecture
@@
The network consists of a contraction path (on the left), an expansion path (on the right), and a Projection-Reprojection (PRP) unit. The contraction path includes four downsampling layers, each consisting of a 3D convolutional layer followed by a max-pooling layer. The convolutional layers apply LeakyReLU activation and progressively increase the number of feature channels, while the pooling layers reduce the data dimensionality. The contraction path ends with a connection to the PRP unit.
@@
The expansion path mirrors the contraction path, consisting of four upsampling layers and a single convolutional layer. Each upsampling layer uses a 3D transposed convolution with ReLU activation. Feature maps from the expansion path are combined with those from the contraction path at the same resolution through skip connections before being passed through the upsampling layers. These layers reduce the number of feature channels while increasing spatial resolution. The final upsampling layer produces a 64-channel output, then refined by a output convolution layer to smooth and integrate information across all channels, ensuring a more refined reconstruction.
@@
II-D2 Projection-Reprojection unit
The PRP unit inserted between the encoder and decoder is specially designed for our task. The differences between data modalities cause the encoder and decoder to process semantically distinct information at many low-level layers[30]. To ensure that the network learns a common representation space between the two modalities, it is essential to design an encoding unit that is both expressive and resistant to overfitting.
Therefore, inspired by [30], we design the PRP unit. The structure is detailed in Figure 4. The projection part consists of two fully connected layers. The first fully connected layer, activated by LeakyReLU, compresses the input feature map into a low-dimensional vector. The second fully connected layer eliminates the activation and further compresses the feature’s dimensionality to produce the latent representation $z$. The reprojection part reverses the projection process by symmetrically expanding the feature vector and reshaping it back to the original feature map size. The only difference is that the final fully connected layer uses ReLU activation to ensure consistency in the data dynamics within the network.
@@ II-D2 Projection-Reprojection unit
II-E Differentiable volume rendering
Figure 5: Rendering diagram.
To supervise the 3D reconstruction process using 2D optical images, a projection transformation technique is required to map volumetric 3D data onto a 2D image plane, thus enabling quantitative error evaluation between the two domains. However, to integrate this projection module into neural network training, the transformation process must remain differentiable to ensure gradient backpropagation of training errors. In CMAR-Net, we use differentiable volume rendering[31] to achieve this critical mapping.
Rendering refers to the computational process that converts mathematical representations of 3D scenes (e.g., geometric models, material properties, lighting conditions, and camera parameters) into 2D images. Differentiable volume rendering is a specialized variant for rendering 3D volumetric data(e.g. density fields). Its core principle involves integrating density and color values along light rays to establish a differentiable mapping from the 3D scene to the 2D image.